OpenStudy (anonymous):

evaluate t an(cos^-1(-15/17))

OpenStudy (anonymous):

Let \(\large y=cos^{-1}(\frac{-15}{17}) \) and now your problem is basically asking to evaluate tan y = ??? So \(\large cosy=\frac{-15}{17} \)

OpenStudy (anonymous):

Now make a right triangle according to cosy = -15/17. Can you find the missing side here?: